OpenStudy (anonymous):

Area under a curve stuff. Find the area bounded by the curves y^2 = 2x + 6 and x = y + 1. Your work must include an integral in one variable.

OpenStudy (anonymous):

so far i've found that the curves in terms of x are y = sqrt(2x + 6) and y = x-1

OpenStudy (anonymous):

@Hero do you know how to do these ?

OpenStudy (anonymous):

oh okay thanks for stopping by either way :)

OpenStudy (anonymous):

Graph it and find the bounded region in order to set up the integral.

OpenStudy (anonymous):

The way you got the y by itself is correct.

OpenStudy (anonymous):

yep i already graphed it

OpenStudy (anonymous):

which function is on top of the bounded region?

OpenStudy (anonymous):

um the y = sqrt(2x + 6) one right?

OpenStudy (anonymous):

Does the problem also say bounded by x=0?

OpenStudy (anonymous):

no copyed and pasted exactly what was given to me .

OpenStudy (anonymous):

i'd assume it is a given though right?

OpenStudy (anonymous):

It should be or the answer is infinite from just looking at the graph. So the bottom function is y=x-1 \[\int\limits_{0}^{5}\sqrt{2x+6}-(x-1)dx\]

OpenStudy (anonymous):

since x-1 is cut off by the x=0, the square root of (2x+6) remains you add the first integral I gave you to this \[\int\limits_{-3}^{0}\sqrt{2x+6}\]

OpenStudy (anonymous):

Know how I got that?

OpenStudy (anonymous):

yes , why dont you just use the \[\int\limits_{-3}^{5}\sqrt{2x+6}-(x-1)\]

OpenStudy (anonymous):

Because the x=0 cuts the x-1 out. You can see that the bounded region does not touch the x-1 from x=-3 to x=0 Do you see the graph?

OpenStudy (anonymous):

yes i do  is this what that integral would find?